Kevin determined that he could ski at the rate of 6 m in 1/5 of a second. How many meters could Kevin ski in 12 seconds?
Answer by rfer(16322)

6/0.2=x/12
0.2x=72
x=72/0.2
x=360 meters
------------
6/(1/5)=x/12
1/5x=72
x=72/(1/5)
x=72*5/1
x=360 meters